# Perception of Visual Content: Differences Between Humans and Foundation Models

### v1

@@ Conclusion
Answering the RQs. We first look at RQ1: “How similar are human-generated and ML-generated annotations?”. It can be concluded that from a low-level perspective, i.e., what types of word occur, sentence structures, the ML Objects and ML Captions are similar to each other, but are relatively different to the human labels. This is expected due to underlying cognitive processes during the annotation stage. However, when looking at them from a high-level point of view, despite the apparent low similarity, they are alike in that they exhibit consistent similarity or dissimilarity regardless of the geographical region of the images, indicating lack of regional bias. This led to RQ2: “How do different combinations of annotations affect fairness in ML predictive models?” In the case of region classification, it was revealed that the human labels resulted in the best overall-performing model, as well as the most balanced performance across individual classes. It is additionally apparent that each region obtained best classification performance with models that utilised annotation combinations that included human labels. Similarly, for income regression, the model that has the best overall performance also has the most balanced performance across regions. However, in this case, the combination of ML Objects and ML Captions is what results in this enhanced performance. A similar trend continues in that the best performance for individual classes was obtained with models that contained ML Objects in its annotations. Considering these findings, it can be deduced that there is no particular combination of annotations that result in best overall and most balanced performance as that is highly dependent on the quality of the images and how well the human annotator or machine is able to select the best discriminative features. Additionally, due to imbalanced distributions of income across data samples and low number of samples for certain regions, the regression model is prone to underestimation or overestimation of income for certain regions. Moreover, our analysis suggests that it may be plausible that models with the best overall performance tend to also be more balanced on the class-level, and therefore, fairer. However, further studies need to be done to confirm this.
@@
Implications. These findings have several implications. First, despite their apparent differences, humans and machines may be more alike in their bias, or lack thereof, in the context of visual content. Their consistent similarity and dissimilarity of annotations across regions implies that their level of fairness is comparable. This insight could steer researchers towards further advancing ML models to avoid as much bias as possible. Second, achieving a more balanced performance on the region level can help reduce biases in ML models. This would result in models that not only give accurate outcomes, but ones that are also fair. This would be highly beneficial in real-world scenarios where the model’s outcomes could disproportionately affect certain demographic groups. Third, due to machine annotations (ML Objects and ML Captions) performing better and fairer at the regression task, tasks in other domains could potentially benefit from using machine annotations as well. However, this could result in a potential negative societal impact in that humans become overly reliant on machines, which may cause them to miss possible biases in the models that originated from the training data, miss errors due to assuming the superior performance of the ML models, or even cause difficulties in identifying accountability if the ML models’ prediction errors result in significant repercussions. This could also lead to potential misuse of the technology as it is possible that one intentionally uses a model that is potentially biased without attempting to address the fairness problems.
@@ Acknowledgements
ChatGPT and Claude AI were utilised to assist with brainstorming and to generate sections of this Work, including alternative phrasings and parts of code used for the experiments.

### v2

@@ Conclusion
Answering the RQs. We first look at RQ1: “How similar are human-generated and ML-generated annotations?”. It can be concluded that from a low-level perspective, i.e., what types of word occur, sentence structures, the ML Captions and Human Labels are similar to each other, but ML Objects are relatively different to either of them. This is expected due to similarities in underlying cognitive processes between ML models and human brains. However, when looking at them from a high-level point of view, despite the apparent low similarity between ML Objects and the other annotations, all of them are alike in that they exhibit consistent similarity or dissimilarity, even when considering the spread based on region, where all of them have the same trends, with Europe clustering together and showing a strong agreement across all annotation types. This led to RQ2: “How do different combinations of annotations affect fairness in ML predictive models?” In the case of region classification, it was revealed that different annotation combinations work best for different geographical regions and image categories. This points to the conclusions that all annotations are important, and machine-generated annotations cannot just replace human-generated ones. It is additionally apparent that each region obtained best classification performance with models that utilised annotation combinations that included ML Captions. In the case of income regression, the combination of ML Objects and ML Captions is what results in enhanced performance. Unlike the region classification, the best performance for individual classes was obtained from different annotation models. Considering these findings, it can be concluded that there is no particular combination of annotations that result in best region-based performance as that is highly dependent on the quality of the images and how well the human annotator or machine is able to select the best discriminative features.
Implications. These findings have several implications. First, despite their apparent differences, humans and machines may be more alike in their bias, or lack thereof, in the context of visual content. Their consistent similarity and dissimilarity of annotations across regions implies that their level of bias is comparable. This insight could steer researchers towards further advancing ML models to avoid as much bias as possible.
@@
Second, due to machine annotations (ML Captions only and ML Objects-ML Captions) performing better overall at both the classification and regression tasks, tasks in other domains could potentially benefit from using machine annotations as well. However, this could result in a potential negative societal impact in that humans become overly reliant on machines, which may cause them to miss possible biases in the models that originated from the training data, miss errors due to assuming the superior performance of the ML models, or even cause difficulties in identifying accountability if the ML models’ prediction errors result in significant repercussions. This could also lead to potential misuse of the technology as it is possible that one intentionally uses a model that is potentially biased without attempting to address the underlying issues.
@@ Acknowledgements
This work is partially supported by the Swiss National Science Foundation (SNSF) under contract number CRSII5_205975 and by an Australian Research Council (ARC) Future Fellowship Project (Grant No. FT240100022). ChatGPT and Claude AI were utilised to assist with brainstorming and to generate sections of this Work, including alternative phrasings and parts of code used for the experiments.